OpenStudy (anonymous):

If f(x) = x + 4, find f(3).

OpenStudy (saifoo.khan):

f(x) = x + 4 f(3) = 3 + 4 f(3) = 7

OpenStudy (anonymous):

Thank you :)

OpenStudy (saifoo.khan):

No problem! :)

OpenStudy (anonymous):

PRECALCULUS PROBLEM:the arithmetic mean of two positive numbers exceeds their geometric mean by 50. by how much does the square root of the larger exeed the square root of the smaller? hint: consider the factored form of x^2 -2xy +y^2 when appropriate

OpenStudy (saifoo.khan):

Sorry man! if i knew it, i would have really helped. :/

OpenStudy (anonymous):

it's okay thanks anywayy